Two alpha particles (helium nuclei), each consisting of two protons and two neutrons, have an electrical potential energy of 6 × 10−19 J . What is the distance between these particles at this time? The Coulomb constant is 8.98755 × 109 N · m^2 /C^2 , the elemental charge is 1.6021 × 10−19 C, and the acceleration due to gravity is 9.8 m/s^2 . Answer in units of m.
Answer #1

q1 = q2 = 2e = 2(1.6021*10^-19)

= 3.2*10^-19 C

Formula for the electric potential is

PE = k(q1*q2) /r

k(q1*q2) / PE = r

(8.98755*10^9)(3.3*10^-19)(3.2*10^-19)/(6*10^-19) = r

r = 1.53*10^-9 m